# Exploring the Boundaries of On-Device Inference: When Tiny Falls Short, Go Hierarchical

### v1

@@ 5.2. Results
It can be seen that in Table 6, for each model and device, EE-HI outperforms the base model with the HI strategy. Reductions of up to $60$% in both latency and energy consumption are observed for ResNet-56 on Raspberry Pi. Fig. 7 provides a more in-depth comparison. For each of the base models, the performance of all the tested strategies is depicted: base model, base model with HI using binary LR (HI), on-device only early exit (EE), EE-HI using binary LR (EE-HI), and remote inference only with a larger ML model (ViT-H/14). For the two early exit strategies, several threshold alternatives are presented, starting from 0.5 and monotonically increasing until a threshold of 1 is reached. Higher thresholds lead to fewer samples being accepted early, resulting in corresponding increments in both accuracy and latency. Consequently, a trade-off between latency/energy and accuracy can be observed. The point where the dotted lines intersect the QoS requirement limit yields the optimal threshold presented in Table 6.
On-device-only EE models (orange dots) fail to achieve the QoS requirement, regardless of the selected threshold. Furthermore, they also fail to reach the base model accuracy except for AlexNet. However, with the appropriate threshold selection, they can lead to notable improvements in terms of latency and energy with reasonable accuracy.
The more branches that are introduced in the model, the more flexibility there is in terms of the accuracy-latency trade-off. This can be observed for ResNet-56 implementations in Fig. 7. The dark-colored lines depict all the possible configurations attainable once the first threshold is fixed to $\theta_{1}=0.81$. However, to illustrate the concept, two light-colored lines are presented. These are obtained by fixing $\theta_{1}=0.95$ and varying $\theta_{2}$. Although they do not provide the optimal configuration for the given QoS, they can reach much higher accuracy while still benefiting from latency reductions. Consequently, EE-HI emerges as the best strategy under the given circumstances, benefiting from the accuracy boosts of HI and the acceleration of early exit models.
6. Conclusion and Future Work
@@
In this work, we have empirically demonstrated that the HI system complements the on-device inference by providing up to $73$% reduction in latency and $77$% reduction in on-device energy consumption in the studied scenarios. Additionally, we have designed a hybrid EE-HI system that features early exits within the HI framework. This approach allows us to achieve up to 60% reduction in latency and energy consumption compared to HI alone for CIFAR-10 image classification on Raspberry Pi and Jetson Orin Nano. Despite these advancements, using HI remains inferior to remote inference in certain scenarios where no on-device models are small enough to meet latency and energy QoS requirements while maintaining reasonable accuracy. This highlights the need for designing DL models with a better accuracy-to-size ratio than current SOTA on-device models.
@@
Our work opens several new research directions. We used the BranchyNet design approach (Teerapittayanon et al., 2016) for early exit technique. Although pioneering, BranchyNet is not the most sophisticated approach. We aim to explore novel techniques like self-distillation (Zhang et al., 2021), which show significant improvement in early exit implementation by reducing latency and energy consumption and improving accuracy compared to the base model. We also plan to investigate how this concept could identify complex samples earlier in the network, further accelerating system inference. While binary LR is an effective HI offloading decision algorithm, it does result in some false positives (correct local inference samples that are offloaded) and false negatives (incorrect local inference samples that are not offloaded). In our HI systems, the binary LR achieved an F1 score of 0.86 for ResNet-8 on CIFAR-10 and an F1 score of 0.83 for ResNet-18 on ImageNet-1K. We aim to explore other HI offloading decision algorithms that further reduce false positives and false negatives, thereby improving the efficiency of HI systems. Additionally, developing local ML models with a high accuracy-to-size ratio will ensure that these models achieve the stringent QoS requirements with HI while maintaining lower latency and energy consumption. Finally, our experiments used a static setup with a dedicated server and stable WiFi access, resulting in consistent offloading times. Studying the effect of device mobility and network interference on the offloading times and the efficiency of HI systems is left for future work.
Acknowledgement
This work is supported by
@@ Acknowledgement
6.1. Additional measurements
In Table 7, we present the accuracy and latency measurements taken for Arduino Nano and ESP-32 for CIFAR-10 dataset. As mentioned before, these MCUs possess very limited computational capabilities making it hard to satisfy latency requirements for even moderately complex tasks. The average latency for a single on-device inference is approximately $6.5$ times slower than that of remote inference for the Arduino Nano, and nearly $60$ times slower for the ESP-32. It should be noted that remote inference is significantly faster on the ESP-32 because it uses Wi-Fi as its communication protocol, as opposed to Bluetooth used by the Arduino Nano. Therefore, the preferred solution would be to offload the inference task to a powerful edge server and receive the corresponding prediction. This offloading process is labeled as ”Remote” in Table 7. The term ”HI” refers to the entire system, encompassing both on-device inference and remote inference for a specified subset of images, denoted as ”offload,” which indicates the percentage of total images that are offloaded for more detailed inference.
Table 7. ResNet-8 on CIFAR-10 results for Arduino Nano and ESP-32.
@@ 6.1. Additional measurements
In Table 8, we present the latency and accuracy measurements for the base models on the Jetson Orin Nano with different precision levels. The entire test set for CIFAR-10 and the entire validation set for ImageNet-1k have been used for accuracy computation. For latency measurement, the values are averaged over $10000$ images, that are randomly selected. This table reinforces the effects of quantization from a precise measurement standpoint. We can observe that as models are further quantized, the average latency requirements for a single inference reduce gradually without significantly impacting accuracy. Another key observation is that quantization strategies are more beneficial for larger models compared to smaller ones. For example, with the CIFAR-10 dataset, full integer quantization results in a $19$% reduction in latency for ResNet-8 compared to full precision. In contrast, ResNet-56 and AlexNet experience $42$% and $53$% reduction, respectively. For the ImageNet-1K dataset, however, all three models exhibit similar gains in latency reduction, each achieving around $69$%.
Table 8. Latency and accuracy for the base models on the Jetson Orin Nano with different precision levels.

### v2

@@ VII-B Results
A more in-depth comparison is presented in Fig. 8, where the performance of all tested strategies is depicted: base model, base model with HI using binary LR (HI), EE (base model with early exit), EE-HI using binary LR (EE-HI), and remote inference with a larger DL model (ViT-H/14). For the EE and EE-HI strategies, we vary the confidence thresholds from $0.5$ to $1$, incrementally increasing the threshold. Higher thresholds result in fewer samples being accepted early, leading to improved accuracy but increased latency. This highlights the trade-off between latency/energy and accuracy. The optimal thresholds, where the dotted lines intersect the QoS requirement limit, are provided in Table IX for EE-HI.
Note that EE models (orange dots) fail to meet the accuracy QoS requirement, regardless of the selected threshold. Additionally, they achieve lower accuracy than the base models (black cross marker), except for AlexNet. However, with appropriate threshold selection, EE models can provide significant improvements in latency and energy consumption while maintaining reasonable accuracy. Introducing more branches in the model increases flexibility in the accuracy-latency trade-off, as demonstrated by the ResNet-56 implementation in Fig. 8. The dark-colored lines in the figure represent all possible configurations attainable when the first threshold is fixed at $\theta_{1}=0.81$ (with $\theta_{2}$ increasing monotonically). To further illustrate the concept, two light-colored lines are included, showing configurations obtained by fixing $\theta_{1}=0.95$ and varying $\theta_{2}$. While these configurations do not achieve the optimal QoS, they demonstrate the ability to achieve higher accuracy while still benefiting from latency reductions. As a result, EE-HI emerges as the best strategy across all the considered scenarios. It combines the accuracy improvements of HI with the reduced latency and energy benefits of EE models. Moreover, its flexibility enables the development of systems tailored to specific tasks and QoS requirements.
VIII Conclusion and Future Work
@@
In this work, we presented processing times, energy consumption, and accuracy measurements for inference using AlexNet and ResNet variants on five devices: Arduino Nano, ESP32, Coral Micro, Raspberry Pi, and Jetson Orin, across MNIST, CIFAR-10, and ImageNet image classification tasks. Using these measurements, we compared the performance of on-device inference, remote inference, and HI approaches. Unlike existing works, we propose that a HI system with a smaller on-device model is more effective for meeting QoS requirements than a HI system using a SOTA on-device model. Our results show that the HI system complements on-device inference, achieving up to a $73$% reduction in latency and a $77$% reduction in on-device energy consumption in the studied scenarios. Furthermore, we designed a hybrid EE-HI system that integrates EE models into the HI framework. This approach demonstrated up to a $60$% reduction in latency and energy consumption compared to HI alone for CIFAR-10 image classification on Raspberry Pi and Jetson Orin. Despite these gains, we find that in scenarios where no on-device model is small enough to meet latency and energy QoS requirements while maintaining reasonable accuracy, the HI system defaults to offloading all data samples for remote inference. This underscores the need for designing DL models with a better accuracy-to-size ratio than current state-of-the-art (SOTA) on-device models.
@@
Our work opens several new research directions. For the early exit technique, we employed the BranchyNet design approach [24]. We aim to explore novel techniques like self-distillation [51], which have shown significant improvements in early exit implementation by reducing latency and energy consumption while improving accuracy compared to the base model. We also aim to explore more efficient optimization techniques to streamline threshold selection in EE-HI. While binary LR is an effective HI offloading decision algorithm, it produces false positives (incorrect local inference samples that are not offloaded) and false negatives (correct local inference samples that are offloaded). In our HI systems, binary LR achieved an F1 score of 0.86 for ResNet-8 on CIFAR-10 and 0.83 for ResNet-18 on ImageNet-1K. To further enhance the efficiency of HI systems, we plan to investigate alternative offloading decision algorithms that reduce false positives and false negatives. Additionally, developing local ML models with a high accuracy-to-size ratio will enable these models to meet stringent QoS requirements with HI while maintaining low latency and energy consumption. Finally, our experiments were conducted in a static setup with a dedicated server and stable WiFi access, resulting in consistent offloading times. Future work will focus on studying the impact of device mobility and network interference on offloading times and the overall efficiency of HI systems.
@@
Additional measurements
In Table X, we present the accuracy and latency measurements taken for Arduino Nano and ESP-32 for CIFAR-10 dataset. As mentioned before, these MCUs possess very limited computational capabilities making it hard to satisfy latency requirements for even moderately complex tasks. The average latency for a single on-device inference is approximately $6.5$ times slower than that of remote inference for the Arduino Nano, and nearly $60$ times slower for the ESP-32. It should be noted that remote inference is significantly faster on the ESP-32 because it uses Wi-Fi as its communication protocol, as opposed to Bluetooth used by the Arduino Nano. Therefore, the preferred solution would be to offload the inference task to a powerful edge server and receive the corresponding prediction. This offloading process is labeled as “Remote” in Table X. The term “HI” refers to the entire system, encompassing both on-device inference and remote inference for a specified subset of images, denoted as “offload,” which indicates the percentage of total images that are offloaded for more detailed inference.
TABLE X: ResNet-8 on CIFAR-10 results for Arduino Nano and ESP-32.
@@
In Table XI, we present the latency and accuracy measurements for the base models on the Jetson Orin Nano with different precision levels. The entire test set for CIFAR-10 and the entire validation set for ImageNet-1k have been used for accuracy computation. For latency measurement, the values are averaged over $10000$ images, that are randomly selected. This table reinforces the effects of quantization from a precise measurement standpoint. We can observe that as models are further quantized, the average latency requirements for a single inference reduce gradually without significantly impacting accuracy. Another key observation is that quantization strategies are more beneficial for larger models compared to smaller ones. For example, with the CIFAR-10 dataset, full integer quantization results in a $19$% reduction in latency for ResNet-8 compared to full precision. In contrast, ResNet-56 and AlexNet experience $42$% and $53$% reduction, respectively. For the ImageNet-1K dataset, however, all three models exhibit similar gains in latency reduction, each achieving around $69$%.
TABLE XI: Latency and accuracy for the base models on the Jetson Orin Nano with different precision levels.